### v1

# Data Diversity Matters for Robust Instruction Tuning

Abstract
Instruction tuning has emerged as a key step in aligning large language models. One of the central challenges of instruction tuning is dataset selection, as the composition of the instruction tuning dataset can significantly impact downstream performance. In particular, researchers have hypothesized that dataset diversity and dataset quality are important indicators of downstream performance. However, it is not clear how to automatically select high quality and diverse data or how exactly quality and diversity affect instruction following ability. To resolve these issues, we propose a new algorithm, Quality-Diversity Instruction Tuning (QDIT). QDIT provides a principled algorithm to control dataset diversity and quality, allowing us to conduct an in depth study on the effect of diversity and quality on instruction tuning performance. From this study we draw two key insights (1) there is a natural tradeoff between dataset diversity and quality and (2) increasing dataset diversity significantly improves the worst case instruction following performace, therefore improving robustness. We validate the performance of QDIT on several large scale instruction tuning datasets, where we find it can improve worst case performance by 18% while maintaining or improving average performance compared to quality driven baselines.
1 Introduction
Large pre-trained language models have demonstrated a remarkable ability to perform a wide range of natural language processing tasks (Vaswani et al., 2017; Devlin et al., 2018; Radford et al., 2019; He et al., 2020; Brown et al., 2020). Although these models are powerful, pre-trained models such as GPT-3 can be quite difficult to work with and often do not follow user instructions (Brown et al., 2020). To make language models more user-friendly and to unlock instruction-following capabilities, researchers have turned to instruction tuning, in which language models are trained to follow instructions on a small set of example instruction-response pairs (Mishra et al., 2021; Wei et al., 2021; Sanh et al., 2021; Wang et al., 2022a). Instruction tuning has become extremely popular, allowing researchers to train powerful and aligned language models with a relatively small budget (Taori et al., 2023; Chiang et al., 2023; Xu et al., 2023).
One of the main challenges of instruction tuning is selecting the instruction data. Recent works have found that the composition of the instruction-tuning dataset significantly impacts the instruction following ability of the trained model, with most works identifying two qualities that an instruction tuning dataset should have: it should be high quality (Chen et al., 2023; Peng et al., 2023) and it should cover a wide range of tasks (i.e. be diverse) (Wei et al., 2021; Zhou et al., 2023; Gudibande et al., 2023). However, as of yet there has been little study of how these two factors (quality and diversity) affect on one another and how researchers can automatically select high quality and diverse datasets. Instead, current work relies on manual data curation (Zhou et al., 2023) or optimize one quantity only (Chen et al., 2023; Peng et al., 2023; Wang et al., 2022b).
To bridge this gap we introduce a new algorithm, Quality-Diversity Instruction Tuning (QDIT), to select diverse and high quality instruction tuning datasets. QDIT is inspired by submodular optimization (Fujishige, 2005) and works by iteratively selecting the datapoint that will improve the joint quality-diversity score the most. QDIT provides an effective way to control the diversity and quality of the instruction tuning dataset, allowing us to conduct an in-depth study of diversity and quality in instruction tuning. From this study we identify two key findings: (1) there is an inherent tradeoff between dataset diversity and dataset quality and (2) improving dataset diversity primarily improves the worst and average case instruction following ability, while not affecting best case instruction following ability much. Based on these finding, we are able to use QDIT to optimize the quality-diversity tradeoff, improving worst case performance while maintaining or improving best case and average performance for robust instruction following. We extensively validate the QDIT algorithm with experiments on four large scale instruction tuning datasets.
The rest of this article is organized as follows: Section 2 covers related literature, Section 3 presents the QDIT algorithm, Section 4 contains our experimental results, and Section 5 concludes our work.
2 Related Work
Given the widespread availability of large instruction tuning datasets (Ding et al., 2023; Zheng et al., 2023a), a new question has emerged: How should we select a training dataset from existing ones? Here data selection is necessary as training on large scale instruction tuning datasets will have a large cost and result in deterioration of model capabilities (Ouyang et al., 2022; Zhou et al., 2023; Chen et al., 2023). Several strategies have been proposed to answer this question, mainly focusing on improving dataset quality. Li et al. (2023) propose to select instructions based on difficulty. Chen et al. (2023) propose Alpagasus, in which a large language model such as ChatGPT is employed to select the highest quality (instruction, response) pairs from a large dataset. Peng et al. (2023) also seek to improve data quality, and do so by rewriting answers to instructions with GPT-4. We seek to improve upon these algorithms by including both diversity and quality in the selection criteria.
Our work is also related to several works that use submodular optimization to increase dataset diversity in natural language processing (Kumar et al., 2019; Kirchhoff and Bilmes, 2014). However, to the best of our knowledge we are the first to include both data quality and diversity in the selection procedure.
3 Method
Before presenting QDIT, we discuss how to quantify and optimize dataset diversity.
3.1 Quantifying Dataset Diversity and Quality
Given a set $A\subseteq V$, a natural way to measure the diversity of the set $A$ with respect to $V$ is by the facility location function (Cornuéjols et al., 1983)
$d(A)=\sum_{v\in V}\textrm{max}_{a\in A}sim(a,v)$ (1)
Intuitively, we can see that a set A that has a high diversity score $d(A)$ will have an element $a\in A$ close to each element $v\in V$ and will therefore be representative of the set $V$.
In order to optimize the facility location function, researchers typically use the greedy data selection algorithm (Nemhauser et al., 1978), where the datapoint that most increases the diversity score of the current subset is selected at each iteration. The greedy algorithm is shown in Algorithm 1, where $d(a|A)=d(a\cup A)-d(A)$. We remark that this algorithm achieves the best possible approximation ratio (in the worst case) that a polynomial time algorithm can achieve unless $P=NP$ (Nemhauser et al., 1978).
To measure dataset quality, we follow prior works and measure the quality of each (instruction, response) pair using a large language model such as ChatGPT (Chen et al., 2023) or measure the quality of each datapoint using a reward model trained on large amounts of human preference data (Ouyang et al., 2022; Bai et al., 2022). We refer to such a function with $q(\cdot)$, and measure a dataset’s overall quality by averaging the quality score of each datapoint.
3.2 Quality-Diversity Instruction-Tuning
In order to select data based on both quality and diversity, we alter the greedy algorithm so that instead of selecting based on diversity alone at each iteration, we select data based on a linear combination of quality and diversity $(1-\alpha)d(a)+\alpha q(a)$. Here $\alpha$ is a hyperparameter that can be used to control the diversity tradeoff. We remark that when $\alpha=1$, QDIT becomes the quality driven selection algorithm proposed in Chen et al. (2023) and when $\alpha=0$ QDIT becomes the classical greedy algorithm (Nemhauser et al., 1978). QDIT will select points that most increase the joint diversity-quality score of the data set, therefore satisfying our goal of building a diverse and high quality dataset. The QDIT algorithm can be found in Algorithm 2. In QDIT, we use the cosine similarity of instruction embeddings as the similarity function in (1), where the instruction embeddings are computed with sentence transformers (Reimers and Gurevych, 2019). See appendix F for more details.
Algorithm 1 Greedy Data Selection
1:$K$triangleright Desired subset size
2:$A\leftarrow\emptyset$
3:$R\leftarrow V$
4:$N\leftarrow 0$
5:while $N<K$ do
6:     $a\leftarrow\textrm{argmax}_{a\in R}d(a|A)$
7:     $A\leftarrow A\cup\{a\}$
8:     $R\leftarrow R\setminus a$
9:     $N\leftarrow N+1$
10:end while
Algorithm 2 QDIT Data Selection
1:$K,\alpha$
2:$A\leftarrow\emptyset$
3:$R\leftarrow V$
4:$N\leftarrow 0$
5:while $N<K$ do
6:     $a\leftarrow\textrm{argmax}_{a\in R}(1-\alpha)*d(a|A)+\alpha*q(a)$
7:     $A\leftarrow A\cup\{a\}$
8:     $R\leftarrow R\setminus a$
9:     $N\leftarrow N+1$
10:end while
Computational Complexity of QDIT. Finding $a\in A$ that maximizes the quality-diversity score at each time step has complexity $\mathcal{O}(|V|^{3})$, leading to a total complexity of $\mathcal{O}(|V|^{3}K)$ for QDIT based data selection. This is problematic, since for some datasets $|V|$ is greater than 1 million. We take several approaches to reduce data selection cost: (1) parallelization on GPUs, (2) employing the lazy greedy selection algorithm of Minoux (2005), and (3) sub-sampling according to Mirzasoleiman et al. (2015). These methods can be employed simultaneously, allowing us to select from millions of datapoints within a few hours. We provide more details in Appendix D, where we find that data selection has minimal cost when compared to training.
4 Experiments
We first analyze how the QDIT algorithm affects dataset quality and diversity and then present our main results and analysis.
4.1 Preliminary Analysis
To verify that QDIT can indeed control the quality and diversity of the selected dataset, we first investigate whether the facility location function is aligned with our intuitive understanding of diversity. Next, we study how the $\alpha$ parameter affects the dataset’s quality and diversity. For this analysis, we apply the QDIT algorithm to the Alpaca dataset (Taori et al., 2023).
To qualitatively verify that the facility location function is aligned with dataset diversity, we use the Berkeley Neural Parser (Kitaev and Klein, 2018; Kitaev et al., 2019) to extract the root verb and first direct noun from each instruction in the Alpaca dataset. We then plot the distribution of verb-noun pairs in Figure 1 for random, quality driven, and QDIT data selection. From Figure 1, we observe that selecting based on the facility location function indeed improves the dataset diversity compared to random selection, while selecting based on quality alone decreases dataset diversity.
(a) Random Selection
(b) Quality Driven Selection
(c) QDIT ($\alpha=0$)
Figure 1: Distribution of root verbs and first nouns selected by different algorithms. The dataset size is 3000.
Next we plot how $\alpha$ in QDIT affects dataset quality and diversity in Figure 2. Similar figures for other datasets can be found in Appendix A. From these figures, we can observe that there is a tradeoff between quality and diversity, and that QDIT allows us to smoothly control this tradeoff. Moreover, we observe that QDIT is able to improve diversity without significantly decreasing data quality. Altogether, these results indicate that QDIT is a practical way to control dataset diversity and quality.
(a) 3K Selected Points
(b) 9K Selected Points
Figure 2: Effect of $\alpha$ on dataset quality and diversity. The red line represents a randomly selected dataset.
4.2 Experimental Setup
Now that we can control quality and diversity with QDIT, we seek to study how dataset quality and diversity affect instruction following ability.
$\diamond$ Training Setup. We use QDIT on four instruction tuning datasets: Alpaca 52K (Taori et al., 2023), Ultrachat 1.3M (Ding et al., 2023), LMSYS-Chat 1M (Zheng et al., 2023b), and a combined dataset of Alpaca 52K, Dolly 15K (Conover et al., 2023), and the OIG-small-chip2 dataset (210K). We refer to this dataset as “Mixed 270K”. For each dataset we consider two settings: a small data setting and a large data setting. For our main experiments we follow the training procedure from Taori et al. (2023) and use LLaMA-1 7B as our base model (Touvron et al., 2023). In all settings we use the same number of training epochs, meaning that the training cost is proportional to the number of training instructions. Complete hyperparameter details can be found in Appendix B.
[…]
$\diamond$ Evaluating Robustness. Beyond evaluating average performance on the test dataset, we also evaluate the worst and best case performance of each model. We can evaluate the worst case performance with the HH Score by calculating the average score achieved on the worst 10% of instructions for each model (note that the worst instructions can change depending on the model). Similarly, we can evaluate the worst case performance with AlpacaFarm by measuring the losing rate on the worst 40% (smaller percents reduce the test set size by too much) of instruction categories in the WizardLM evaluation set. Here we use losing rate as a metric because a high losing rate will indicate low robustness. Best case performance is measured in a similar manner with win rate. Measuring best and worst case performance provides more detailed insights into how diversity and quality affect model robustness.
$\diamond$ Baselines. We primarily compare the QDIT algorithm with two baselines: random data selection and quality based selection. We remark that for the Alpaca dataset instruction-response quality is measured according to ChatGPT (using the same setup as Chen et al. (2023)) and for all other datasets quality is measured with the previously mentioned reward model.
We also introduce two variants of the QDIT algorithm that attempt to improve data quality and diversity. The first variant, QDIT:Threshold, first orders the data point by quality, and then iteratively selects instructions that do not have a similarity score greater than $\tau$ with any point included selected subset. This algorithm is different from QDIT in that it prevents the dataset from having similar datapoints, which QDIT allows. The second variant we propose is QDIT-Cluster, in which we first cluster the instructions and then select an equal amount of datapoints from each cluster in a quality-driven manner. Comprehensive details can be found in Appendix C, where we also provide an analysis of how these algorithms select data.
Table 1: Instruction Tuning Results. Random 50K refers to a model trained on a randomly sampled set of 50K points on the corresponding dataset. The top and bottom 40% win and lose rate is computed on the WizardLM evaluation set.
Alpaca 52K Average Performance Worst Case Performance Best Case Performance
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 9K 0.975 - 3.756 0.489 19.23 6.942 21.77
Quality 9K 1.013 - 3.944 0.59 30.38 7.478 31.82
QDIT 9K 1.013 - 3.922 0.573 29.55 7.437 25.71
Random 3K 0.928 - 3.601 0.368 37.35 6.876 26.25
Quality 3K 0.923 - 3.732 0.352 30.63 6.972 20.45
[…]
QDIT:Threshold 3K 0.986 - 3.745 0.452 25.99 6.943 29
QDIT 3K 0.961 - 3.809 0.5 25 7.03 27.06
Mixed 270K Average Performance Worst Case Performance Best Case Performance
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 50K 0.927 1 3.6 0.396 29.06 7.002 25.89
Quality 50K 0.973 1.065 4.295 1.186 34.78 7.746 28.7
QDIT 50K 0.975 1.051 4.33 1.223 35.22 7.761 26.32
Random 10K 0.896 0.985 3.609 0.314 32.69 6.863 24.1
Quality 10K 0.956 1.042 4.342 1.117 33.33 7.952 30
[…]
QDIT:Threshold 10K 0.98 1.06 4.282 1.051 33.96 8.005 30.7
QDIT 10K 0.982 1.063 4.429 1.3 32.65 8.006 26.74
Ultrachat 1.3M Average Performance Worst Case Performance Best Case Performance
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 50K 1.189 1 4.866 0.811 25 8.594 50
Quality 50K 1.249 1.068 5.91 2.116 21.65 9.562 56.25
QDIT 50K 1.26 1.094 5.909 2.185 22.52 9.69 57.45
Random 10K 1.137 0.968 4.891 0.931 27.47 8.756 47.78
Quality 10K 1.228 1.011 5.844 1.846 25.47 9.74 48.15
[…]
QDIT:Threshold 10K 1.242 1.037 5.734 1.937 19.64 9.64 57.73
QDIT 10K 1.243 1.037 5.849 2.044 23.89 9.619 51.19
LMSYS 1M Average Performance Worst Case Performance Best Case Performance
Winning Score vs. Alpaca 52K normal-$\uparrow$ Winning Score vs. Random 50K normal-$\uparrow$ HH Score Mean normal-$\uparrow$ Bottom 10% HH Score normal-$\uparrow$ Bottom 40% Lose Rate normal-$\downarrow$ Top 10% HH Scorenormal-$\uparrow$ Top 40% Win Rate normal-$\uparrow$
Random 50K 1.093 1 4.791 0.958 32.35 8.629 44.21
Quality 50K 1.22 1.182 5.963 2.133 30 9.361 53.06
QDIT 50K 1.217 1.168 6.074 2.278 34.48 9.484 51.55
Random 10K 1.108 1.002 4.906 0.981 26.21 8.67 44.66
Quality 10K 1.191 1.135 6.139 2.335 31.58 9.69 48.28
[…]
QDIT:Threshold 10K 1.22 1.173 5.978 2.319 26.13 9.709 55.42
QDIT 10K 1.217 1.142 6.078 2.465 25.53 9.712 52.5
4.3 Main Results
The main experimental results can be found in Table 1. We mainly focus our analysis on the low data settings, as we find models trained with QDIT can match or beat the models trained in the high data settings while significantly reducing training cost. Nonetheless, we include the high data setting for reference.
Effect of Quality. Similar to prior works, we find that selecting data based on quality significantly improves average performance, improving the winning score versus Alpaca 52K by 5.41% and the average HH score by 17.14% in low data settings. However, we also find that selecting based on quality alone can hurt worst case performance, increasing the bottom 40% losing rate in two out of four settings compared to random data selection. We hypothesize that this performance drop is due to the fact that quality based selection hurts dataset diversity.
Effect of Diversity. On the other hand, we find that data selection with QDIT is able to achieve both a high average HH score (QDIT improves upon quality driven selection by 2.56% for winning score vs Alpaca 52K and improves average HH score by 1.29% in low data settings) while achieving a much better worst case performance than both random and quality-driven selection. In particular, we find that QDIT improves worst case HH score by 18.68 % and decreases worst case losing rate vs Alpaca 52K by 11.45% when compared to quality-driven data selection. We suspect that QDIT’s more diverse dataset teaches the model to a respond to a wide range of instructions, therefore decreasing the probability that it fails to follow evaluation instructions. We find that QDIT can also maintain best case performance when compared to quality-diven selection, as the best case HH score is improved by 0.12%. From these experiments we conclude that by increasing data diversity while maintaining data quality, QDIT can improve robust instruction following capability.
QDIT Variants. We also observe that the two variants of QDIT can improve average performance compared to random data selection, but they do not necessarily improve worst case performance compared to quality based selection. Our experiments suggest that clustering based data selection is highly sensitive to the performance of the clustering algorithm. For example, clustering may be easy on datasets like Alpaca which are derived from a small set of seed tasks, but difficult for larger and human generated instructions like LMSYS. For threshold based selection, we hypothesize that constraining intra-datapoint similarity is not as important as selecting instructions that cover the entire training set.
4.4 Analysis
Effect of $\alpha$. We plot the effect of $\alpha$ on average and worst case performance in Figure 3. From Figure 3, we can see that QDIT is relatively robust to the value of $\alpha$, with values of $\alpha\in\{0.5,0.7,0.9\}$ typically having the highest worst case and average performance. However, decreasing $\alpha$ by too much (i.e. $\alpha=0.1$) will result in a significant drop in performance, as will increasing $\alpha$ by too much ($\alpha=1$). This highlights the need for a careful tradeoff between dataset quality and diversity.
(a) Alpaca 3K
(b) Mixed 10K
(c) Ultrachat 10K
(d) LMSYS 10K
Figure 3: Effect of $\alpha$ on best case and worst case performance. The red line represents a randomly selected dataset. Worst reward refers to the bottom 10 percent of HH scores.
Benchmark Performance. For a more comprehensive evaluation of QDIT, we follow InstructEval (Chia et al., 2023) and evaluate our model on various benchmark datasets including MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), DROP (Dua et al., 2019), and HumanEval (Chen et al., 2021). Although these benchmarks are not fully aligned with instruction following performance, we find that QDIT uniformly improves MMLU performance and performs best on BBH and DROP in three out of the four scenarios when compared with quality driven selection. These results further confirm the efficacy of QDIT.
Table 2: Evaluation on benchmark datasets. We bold the best result out of quality based selection and QDIT based selection. Details can be found in Appendix G.
Alpaca 9K Mixed 10K
MMLU BBH DROP HumanEval MMLU BBH DROP HumanEval
Random 34.18 31.69 26.84 8.54 32.93 30.92 20.33 10.37
[…]
QDIT 35.64 30.85 26.55 7.93 33.29 32.36 26.19 10.37
Ultrachat 10K LMSYS 10K
MMLU BBH DROP HumanEval MMLU BBH DROP HumanEval
Random 32.12 33.19 26.24 13.41 33.05 32.57 25.06 15.85
[…]
QDIT 35.78 32.66 26.73 13.41 35.9 32.76 26.22 11.58
Different Base Models. We evaluate the performance of QDIT with different base models in Figure 3(a). From Figure 3(a) we find that QDIT can improve both average performance and worst case performance for other base models, supporting our main results. Interestingly we find that the Mistral model performs best according to HH score, but worst according to AlpacaFarm evaluation (Jiang et al., 2023). This highlights the importance of using multiple evaluation strategies.
(a) Different Base Models
(b) Different Data Sizes (Thousands)
Figure 4: (a) Ablation with different base models. The train set is 10K instructions selected from Ultrachat. (b) Ablation with different data sizes. The worst reward is computed as the bottom 10 percent of HH scores, and the winning score is vs. Alpaca 52K.
Data Size. We evaluate the performance of QDIT with different training data sizes in Figure 3(b), where we find that QDIT results in a 3-4 times improvement in data efficiency (and therefore training cost) as measured by robust performance. We also find that increasing the dataset size beyond $50K$ only results in marginal gains of average performance, and can even decrease worst case HH score. We hypothesize this phenomena is due to overfitting.
Case Study. We conduct a case study in Figure 5. We observe that quality based selection does not select an instruction similar to the test instruction and fails to respond correctly. On the other hand, QDIT does train on a similar instruction due to its more diverse training set and correspondingly responds to the instruction correctly. We remark that in practice the distance to the closest training instruction does not always correlate to evaluation performance, as we find embedding distance may not capture important aspects of an instruction such as the format it is in or the desired response format. In the future, we hope to investigate how different similarity metrics impact QDIT and instruction tuning performance.
Figure 5: Spotify playlist case study.
5 Conclusion
In this paper we introduce QDIT, a method for automatically selecting high quality and diverse instruction tuning datasets. QDIT allows us to flexibly tradeoff quality and diversity, improving both average performance and model robustness on a wide variety of datasets. As future lines of research, we would like to investigate how different similarity metrics affect the performance of QDIT and how QDIT performs with different model sizes.
Appendix A Tradeoff between Quality and Diversity: Additional Results
We display additional results on the effect of $\alpha$ on dataset diversity and quality in Figures 6, 7, 8. We again find that there is a tradeoff between quality and diversity, and that $\alpha$ can be used to control this tradeoff.
(a) 1K Selected Points
(b) 5K Selected Points
(c) 10K Selected Points
Figure 6: Effect of $\alpha$ on dataset quality and diversity. The dataset is Mixed 270K.
(a) 1K Selected Points
(b) 5K Selected Points
(c) 10K Selected Points
Figure 7: Effect of $\alpha$ on dataset quality and diversity. The dataset is Ultrachat 1.3M.
(a) 1K Selected Points
(b) 5K Selected Points
(c) 10K Selected Points
Figure 8: Effect of $\alpha$ on dataset quality and diversity. The dataset is LMSYS 1M.
Appendix B Training Details
We display the hyperparameter details in Table 3 and Table 4.
Table 3: General training details. The same hyperaparameter setting is used for every dataset and data selection strategy, following Chen et al. [2023].
Batch Size Learning Rate Epochs Max Length Weight Decay
128 $2\times 10^{-5}$ 3 512 0
Table 4: Selected values of $\alpha$.
Dataset Size Alpaca 52K Mixed 270K Ultrachat 1.3M LMSYS 1M
3K 0.7 - - -
9K 0.7 - - -
10K - 0.7 0.7 0.7
50K - 0.9 0.9 0.7
Appendix C QDIT Variants
In this section we describe and analyze the variants of QDIT.
QDIT:Cluster. In this variant of QDIT, we first cluster all the instructions based on their sentence-transformer embeddings using the $k$-means algorithm, where $k=100$. We then select an equal number of points from each cluster, and points are selected from each cluster based on quality.
QDIT:Threshold. In this variant of QDIT, we first sort the instructions based on quality. We then iteratively remove instructions from the selected dataset if they have a similarity with some other instruction in the dataset greater than a threshold $\tau$. In our experiments we use the cosine similarity as the similarity metric and set $\tau=0.5$.
Analysis. To further analyze why QDIT typically outperforms the two QDIT variants, we plot the selection strategy of each variant on a sample problem in Figure 9. From this figure we can observe the QDIT:cluster will sample equally from across the sample space, leading to some poor quality datapoints being sampled. We also observe that QDIT:Threshold does not necessarily optimize for diversity and may not sample from large portions of the sample space. On the other hand QDIT samples only high quality datapoints from each portion of the sample space, therefore resulting in a high quality and diverse dataset.
(a) QDIT:Cluster
(b) QDIT:Threshold
(c) QDIT
Figure 9: We display the selection strategy of the various QDIT algorithms on a test dataset. Six datapoints are selected and the selected datapoints are circled.
Appendix D Computational Cost of QDIT
We measure the wall time of QDIT as taking approximately 9.42 minutes to select a dataset of size 10000. This experiment was conducted on a single A100 40G GPU. In contrast, training on Ultrachat 10K takes 48 minutes on 8 A100 40G GPUs, meaning that data selection (given the embeddings and reward score) has 2% of the cost of training. Taking into account the inference and evaluation process, our method has an even smaller relative cost. We remark that the reward score and embedding generation can typically be done on a small GPU very quickly, or it can even be done on a CPU for minimal cost.
Appendix E Complete Win-Tie-Lose Results
In this appendix we show the win, ties, and losses achieved versus the reference models as judged by Claude 2. The results can be seen in Figures 10, 11, 12, 13, 14, and 15.
(a) Random 10K vs. Alpaca 52K
(b) Quality 10K vs. Alpaca 52K
(c) QDIT 10K vs. Alpaca 52K
Figure 10: We display the results on Ultrachat as judged by Claude 2. The base model here is LLaMA-2 7B.
(a) Random 10K vs. Alpaca 52K
(b) Quality 10K vs. Alpaca 52K
(c) QDIT 10K vs. Alpaca 52K
Figure 11: We display the results on Ultrachat as judged by Claude 2. The base model here is LLaMA-1 7B.
(a) Random 10K vs. Random 50K
(b) Quality 10K vs. Random 50K
(c) QDIT 10K vs. Random
50K
Figure 12: We display the results on Ultrachat as judged by Claude 2. The base model here is LLaMA-1 7B.
(a) Random 10K vs. Alpaca 52K
(b) Quality 10K vs. Alpaca 52K
(c) QDIT 10K vs. Alpaca 52K
Figure 13: We display the results on LMSYS as judged by Claude 2. The base model here is LLaMA-1 7B.
(a) Random 10K vs. Random
50K
(b) Quality 10K vs. Random
50K
(c) QDIT 10K vs. Random
50K
Figure 14: We display the results on LMSYS as judged by Claude 2. The base model here is LLaMA-1 7B.
(a) Quality 10K vs. Alpaca
52K
(b) QDIT 10K vs. Alpaca
52K
Figure 15: We display the results on Ultrachat as judged by Claude 2. The base model here is Mistral 7B.
Appendix F Similarity Metrics
In QDIT, we use the cosine similarity of instruction embeddings as the similarity function in (1), where the instruction embeddings are computed with sentence transformers [Reimers and Gurevych, 2019]. More specifically, we use the all-mpnet-base-v2 model available from Huggingface.
Appendix G Benchmark Environments
For evaluation on common NLP benchmarks, we follow Chia et al. [2023]. In particular, we conduct five shot evaluation on MMLU, three shot on BBH, three shot on DROP, and zero shot on HumanEval.